# Real-time Estimation of Bound Water Concentration during Lyophilization with Temperature-based State Observers

### v1

@@ Introduction
where $\mathbf{F_{ref}}=\mathbf{F(x_{ref})}$ is the function $\mathbf{F}$ evaluated at $\mathbf{x_{ref}}$, $\mathbf{F^{\prime}_{ref}}=\mathbf{F^{\prime}(x_{ref})}$ is the Jacobian $\mathbf{F^{\prime}}$ evaluated at $\mathbf{x_{ref}}$, and $\mathbf{x_{ref}}$ is the reference state. The average temperature and concentration are used as the reference state, which is constant and uniform. Note that, although there is no real-time concentration measurement available, the average concentration used in this analysis can be obtained via offline measurement during the design and development phase as discussed in the Results section. Alternatively, the literature value could be used, but that would result in a less accurate analysis. With (40), the linearized state observer is
@@
$\dfrac{d}{dt}(\hat{\mathbf{x}}-\mathbf{x})=\mathbf{F^{\prime}_{ref}}(\hat{\mathbf{x}}-\mathbf{x})+\mathbf{L}(\hat{\mathbf{y}}-\mathbf{y}),$ (42)
@@
where $\mathbf{I}_{m}$ is an $m$$\times$$m$ identity matrix and $\mathbf{0}_{m,m}$ is an $m$$\times$$m$ zero matrix. With this definition, it is easy to see that (43) and (44) are identical to (26) and (27). Substituting (43) and (44) into (41) followed by rearranging gives that
$\dfrac{d}{dt}(\hat{\mathbf{x}}-\mathbf{x})=(\mathbf{F^{\prime}_{ref}}+\mathbf{LC})(\hat{\mathbf{x}}-\mathbf{x})-\mathbf{Ln},$ (46)
which is a linear ODE that gives the criterion for observer design, where the eigenvalues of the matrix $\mathbf{F^{\prime}_{ref}}+\mathbf{LC}$ characterize the dynamics of the observer. For the estimation error $\hat{\mathbf{x}}-\mathbf{x}$ to converge to zero, the observer gains $L_{T}$ and $L_{c}$ must be chosen such that the real parts of all eigenvalues of the matrix $\mathbf{F^{\prime}_{ref}}+\mathbf{LC}$ are negative. Besides, the estimation error should decay significantly faster the time scale of the process, i.e., the eigenvalues of the original matrix $\mathbf{F^{\prime}_{ref}}$. Oscillation should also be minimized, which is governed by the imaginary parts of all eigenvalues. This approach allows the systematic selection of the values of $L_{T}$ and $L_{c}$ to specify the dynamics of the observer.
An important parameter of interest that can be extracted from the above analysis is a time constant, denoted as $\tau$, which describes how fast the estimation error reduces to zero. Before discussing the time constant, denote the eigenvalues of the matrix $\mathbf{F^{\prime}_{ref}}+\mathbf{LC}$ from the largest to the smallest as $\lambda_{1},\lambda_{2},\dots,\lambda_{2m}$, in which $\lambda_{1}$ is the largest and $\lambda_{2m}$ is the smallest. As $\mathbf{F^{\prime}_{ref}}+\mathbf{LC}$ is a $2m$$\times$$2m$ matrix, there are $2m$ eigenvalues in total. The simplest choice for approximating the time constant is to consider the slowest eigenvalue:
$\tau=\frac{1}{|\mathrm{Re}(\lambda_{2m})|}.$ (47)
@@
$\tau=\frac{1}{|\mathrm{Re}(\lambda_{m+1})|},$ (48)
which is justified in Supplementary Information by a detailed mathematical analysis on the effects of all eigenvalues. The approximation (48) is sufficiently accurate for all parameter values for the mechanistic model, numerical methods, and state observer used in this work given that the estimated state does not severely oscillate or blow up; i.e., the observer is well designed. Information about the time constant is valuable for gain scheduling when the observer is used to handle strong sensor noise.
The above analysis is based on a linearized version of the observer, and so the result is only an approximation. Nevertheless, this approximation is sufficiently good for practical applications, in particular gain scheduling, as shown in Section 4.4.
@@
Since the mechanistic model is a basis for state observer and control design, this section extensively validates the proposed model with three different datasets from the literature to ensure that our model provides an accurate prediction of the bound water concentration and product temperature.
The first dataset is obtained from the experimental data presented in [15], where the time profile of the total mass of residual water (bound water) during secondary drying was reported. Our mechanistic model can accurately predict the concentration of bound water, with the maximum error is about 0.01 kg water/kg solid (Fig. 3A). The concentration decreases exponentially following the linear driving force model.
The second dataset for model validation is the simulation result obtained from the high-fidelity model proposed by [16]; the model simulates simultaneous heat and mass transfer in 2D. Our model prediction agrees well the result obtained from the high-fidelity model (Fig. 3B), indicating that our simplified model, which simulates the system in 1D and omits mass transfer equations, can still accurately predict the time evolution of the bound water concentration, agreeing with the observation by [33].
The final dataset is obtained from [18], where the time profiles of both residual water and product temperature were reported. Our model can precisely simulate the product temperature (Fig. 3c). The model can also reasonably predict the concentration of bound water despite high uncertainty in the reported measurement.
@@
This section focuses on the simulation-based observer design, with the default parameter values (see all parameter values in Supplementary Information). Practical observer design can be performed by running a series of simulations with different values of observer gains to understand the dynamics of the observer and find gain values that give satisfactory performance. Firstly, consider the observer gain $L_{T}$, which is directly related to the product temperature (Fig. 4A). The estimated average product temperature converges to the true value within 1 h for $L_{T}=-1$$\times$$10^{-6}$ without any oscillation, which is relatively fast given the time scale of 12 h. Some oscillation is observed for $L_{T}$ increased to $5$$\times$$10^{-5}$. The system is highly unstable and the estimated temperature blows up for $L_{T}=1$$\times$$10^{-4}$, indicating that this observer gain is too high and suggesting that the sign of $L_{T}$ is incorrect. Physically, when the estimated temperature is higher than the actual value, the observer term $\mathbf{L_{T}}(\hat{\mathbf{T}}-\mathbf{T})$ should provide negative feedback to the observer to reduce the temperature, so $L_{T}$ should be negative, agreeing with the result from Fig. 4A. With all the analysis in this part, $L_{T}$ is set to $-1$$\times$$10^{-6}$.
Next, consider the observer gain $L_{c}$ (Fig. 4B). This gain is more crucial than $L_{T}$ as it is associated with the concentration of bound water, the most important state to be monitored during secondary drying. The estimated concentration blows up for $L_{c}=-1$$\times$$10^{-7}$, suggesting that the sign of $L_{c}$ is wrong. This behavior is undesirable as it gives unrealistic process values, which could lead to errors or failure in related systems, e.g., control systems. With the positive observer gain $L_{c}=1$$\times$$10^{-8}$, convergence can be observed. However, it takes more than 8 h to achieve the convergence, meaning that the observer reports an incorrect estimated state for most of the process operation. An observer with this sluggish convergence therefore is not beneficial. With a proper gain value, $L_{c}=5$$\times$$10^{-7}$, the estimated concentration converges to the correct value without any oscillation in less than 2 h, which is fast compared to the time scale of 12 h. A slightly slower convergence is observed for the high case, which is reasonable as the initial estimated state is farther from the correct value. When the estimated temperature is higher than the actual value, it physically implies that the estimated desorption rate is too low. Therefore, the observer term $\mathbf{L_{c}}(\hat{\mathbf{T}}-\mathbf{T})$ should provide positive feedback to the observer to increase the estimated concentration, and $L_{c}$ should be positive. Based on this analysis, the observer gain $L_{c}$ is set to $5$$\times$$10^{-7}$.
In Figs. 4AB, the convergence is demonstrated via the average properties as it is intuitive and easy to visualize, although the mechanistic model and state observer also estimate the spatial variation of those properties. In Fig. 4C, the spatiotemporal evolution of the estimation errors is shown for both product temperature and bound water concentration, for $L_{T}=-1$$\times$$10^{-6}$ and $L_{c}=5$$\times$$10^{-7}$. The estimation errors decrease to zero in less than 2 h, showing that the observer can accurately estimate the spatial temperature and concentration, in addition to the average values.
Of course, there are other values for $L_{T}$ and $L_{c}$ that can give satisfactory performance. To provide a systematic study of the choice of observer gains, Fig. 4D plots the convergence times***The convergence time is defined as the time required for the estimation error for concentration is less than 2% of its initial error. for different pairs of $L_{T}$ and $L_{c}$; this plot defines a design space for $L_{T}$ and $L_{c}$ where convergence can be achieved. The plot also highlights the fact that, although the estimation of the concentration is most important, both $L_{c}$ and $L_{T}$ need to be chosen well due to coupling between the species concentration and temperature equations, and a poor choice of value for $L_{T}$ can result in slow convergence. In the later sections of this work, $L_{T}=-1$$\times$$10^{-6}$ and $L_{c}=5$$\times$$10^{-7}$ are used as the default observer gains. This exact same analysis and design procedure can be used for the modified observer, with the final observer gains selected to be $L_{T}=-5$$\times$$10^{-3}$ and $L_{c}=1$$\times$$10^{-4}$.
@@
Another important aspect of the observer is initialization. For the estimated temperature, it is logical to set the initial condition to be equal to the measured temperature. In Fig. 4, the initial estimated temperature is set to be higher than the measured value by 10% to demonstrate the convergence. The initial estimated concentration is, however, unknown as due to no real-time concentration measurement. In Fig. 4, we initialize the observer with the maximum and minimum concentrations (0.6415 kg water/kg solid and 0.0314 kg water/kg solid) reported in the literature [15, 41] to demonstrate the convergence under the worst-case scenarios. In practice, the initial estimated concentration could be set to some more realistic value specific to that system or experiment. This information should be obtained during the design and development phase, which could result in faster convergence. In the later sections of this work, the initial estimated concentration is set to 0.0314 kg water/kg solid and the initial estimated temperature is set to the measured value unless otherwise specified so that the observer performance is analyzed on the same basis.
Figure 4: (A) Convergence of the estimated average product temperature for three different values of $L_{T}$ while $L_{c}$ is kept constant. The initial estimated temperature is set to be higher than the actual value by about 10% to demonstrate convergence. The initial concentration is treated as being known in this plot. (B) Convergence of the estimated average concentration of bound water at three different values of $L_{c}$ while $L_{T}$ is kept constant. The initial estimated concentration is set to the maximum (high) and minimum (low) realistic values reported in the literature, which are about 0.6145 kg water/kg solid and 0.0314 kg water/kg solid, respectively [15, 41]. The initial estimated temperature is set to be higher than the actual value by about 10%. (C) Spatiotemporal evolution of the estimation errors for both temperature and concentration (high) to indicate the convergence at all locations with the selected gains. d, Design space showing the times required for the estimated concentration to converge to the true value at different pairs of the observer gains $L_{T}$ and $L_{c}$. Values of $L_{T}$ and $L_{c}$ outside this design space could lead to severe oscillation or divergence. All simulations are based on the default parameter values.
@@
Figure 5: (A) Convergence of the estimated concentration for different frequency factors $A$. The frequency factor is varied between 7.8$\times$10-5 s-1 [16] and 1.1$\times$10-4 s-1 [14] given that the rate constant is independent on temperature. (B) Convergence of the estimated concentration for different activation energies $E_{a}$. The activation energy is varied between 5,920 J/mol [18] and 13,146 J/mol [42]. (C) Convergence of the estimated concentration for different initial concentrations, $c_{s,0}$. The initial condition is varied between 0.0314 kg water/kg solid [4]) and 0.6415 kg water/kg solid [15]. In all cases, the simulations are terminated when the estimated concentration drops below 0.01 kg water/kg solid. The default observer gains $L_{T}=-1$$\times$$10^{-6}$ and $L_{c}=5$$\times$$10^{-7}$ are used. Other parameters are kept at the default values.
@@
Using the default parameter values, independent normally distributed noise of standard deviation given by $3\sigma=5^{\circ}$ is added to the temperature measurement [40]. The estimated concentration is severely polluted by the measurement noise for the default observer gain $L_{c}=5$$\times$$10^{-7}$ (Fig. 6A). Significant oscillation is observed, in which the observer does not give valuable information when the concentration is small (i.e., after about 6 h). To reduce the noise effect, the observer gain $L_{c}$ is reduced to $2$$\times$$10^{-7}$. In comparison to $L_{c}=5$$\times$$10^{-7}$, the convergence is achieved slightly slower, but the oscillation is much weaker. Reducing the gain $L_{c}$ further to $1$$\times$$10^{-7}$ almost removes the oscillation, but the convergence is also significantly slower.
The above information motivates the use of observer gain scheduling, which is a technique that varies the observer gains with time to ensure satisfactory performance at different operating points, especially for nonlinear systems. In this case, we can initialize the observer with a high gain to achieve fast convergence, and then gradually switch to a low gain after some time to reduce oscillation. Here the switching time is set to $t=4\tau$, where $\tau$ is the time constant derived in Section 3.3. From the definition of the time constant, $t=4\tau$ should be the time when the estimation error is about 98% of the initial value, that is, when the estimated state has nearly converged to the true value. This technique enables fast convergence while having low effects of measurement noise on the state estimate (Fig. 6B).
@@
Note that the above noise level (Fig. 6C) is significantly higher than in actual experiments [40], and so the resulting design has a large safety margin to span the range of noise levels encountered in real temperature sensors. A lower safety margin, with faster convergence to the states, could be achieved by performing the analysis for a noise level set by experimental data for the specific temperature sensor used in the specific equipment.
Figure 6: (A) Convergence and oscillation behavior of the estimated concentration at different observer gains $L_{c}=5$$\times$$10^{-7}$, $2$$\times$$10^{-7}$, and $1$$\times$$10^{-7}$ while $L_{T}$ is fixed at $-1$$\times$$10^{-6}$ in all cases. (B) Gain scheduling technique where $L_{c}$ is set to $5$$\times$$10^{-7}$ for fast convergence at the beginning and then reduced to $1$$\times$$10^{-7}$ at $t=4\tau$ to reduce the effect of measurement noise. (C) Noisy temperature data obtained from adding independent normally distributed noise of standard deviation given by $3\sigma=5^{\circ}$ to the true state. Other parameters kept at their default values.
@@
In the simulation-based observer design, the true and estimated states are simulated simultaneously and continuously. In real systems, measurement data are usually sampled and fed to a state observer in a discrete fashion, i.e., with a fixed sampling time. This section applies the proposed observer to estimate the concentration of bound water in four different experiments from the literature, denoted as Cases A [15], B [1], C [41], and D [4]. The temperature measurement is obtained and fed to the observer every 10 seconds. Parameter values and data preprocessing are described in Supplementary Information.

### v2

@@ Introduction
where $\mathbf{F_{ref}}=\mathbf{F(x_{ref})}$ is the function $\mathbf{F}$ evaluated at $\mathbf{x_{ref}}$, $\mathbf{F^{\prime}_{ref}}=\mathbf{F^{\prime}(x_{ref})}$ is the Jacobian $\mathbf{F^{\prime}}$ evaluated at $\mathbf{x_{ref}}$, and $\mathbf{x_{ref}}$ is the reference state. The average values of temperature and concentration between the initial and final times are used as the reference state, which is constant and uniform.******Although there is no real-time concentration measurement available, the average concentration used in this analysis can be obtained via offline measurement during the design and development phase. Alternatively, a literature value could be used. With (41), the linearized state observer is
@@
$\dfrac{d}{dt}(\hat{\mathbf{x}}-\mathbf{x})=\mathbf{F^{\prime}_{ref}}(\hat{\mathbf{x}}-\mathbf{x})+\mathbf{L}(\hat{\mathbf{y}}-\mathbf{y}).$ (43)
@@
$\mathbf{I}_{m}$ is an $m$$\times$$m$ identity matrix, and $\mathbf{0}_{m,m}$ is an $m$$\times$$m$ zero matrix. With this definition, it is easy to see that (44) and (45) are identical to (27) and (28). Substituting (44) and (45) into (42) followed by rearranging gives that
$\dfrac{d}{dt}\left(\hat{\mathbf{x}}-\mathbf{x}\right)=\left(\mathbf{F^{\prime}_{ref}}+\mathbf{LC}\right)\left(\hat{\mathbf{x}}-\mathbf{x}\right)-\mathbf{Ln},$ (47)
which is a linear ODE that gives the criterion for observer design. The eigenvalues of the matrix $\mathbf{F^{\prime}_{ref}}+\mathbf{LC}$ characterize the dynamics of the observer. For the estimation error $\hat{\mathbf{x}}-\mathbf{x}$ to converge to zero, the observer gains $L_{T}$ and $L_{c}$ must be chosen such that the real parts of all eigenvalues of the matrix $\mathbf{F^{\prime}_{ref}}+\mathbf{LC}$ are negative. Besides, the estimation error should decay significantly faster than the slowest time scale of the process. Oscillation should also be minimized, which is governed by the imaginary parts of all eigenvalues.
The values of the $L_{T}$ and $L_{c}$ in (47) can be systematically selected to give a desired convergence time for the observer dynamics, which can be computed from its slowest time scale $\tau$. Denote the eigenvalues of the matrix $\mathbf{F^{\prime}_{ref}}+\mathbf{LC}$ from the fastest to the slowest as $\lambda_{1},\lambda_{2},\dots,\lambda_{2m}$, in which $\lambda_{1}$ is the fastest and $\lambda_{2m}$ is the slowest. As $\mathbf{F^{\prime}_{ref}}+\mathbf{LC}$ is a $2m$$\times$$2m$ matrix, the matrix has $2m$ eigenvalues in total. The slowest time constant can be estimated from the slowest eigenvalue:
$\tau=\frac{1}{\left|\mathrm{Re}(\lambda_{2m})\right|}.$ (48)
@@
$\tau=\frac{1}{\left|\mathrm{Re}(\lambda_{m+1})\right|},$ (49)
[…]
@@
In this work, the mechanistic model is used in various applications with different operating conditions, which requires a number of input parameters. The default parameter values are given in Table 1. Parameter values different from those reported in Table 1 are stated explicitly in each specific section or case study.
Table 1: Default parameters for simulations.
Parameter Value Unit Reference
$\rho$ 215 kg/m3 [6]
$\rho_{d}$ 212.21 kg/m3 [6]
[…]
$\Delta H_{s}$ $2.68$$\times$$10^{6}$ J/kg [6]
$E_{a}$ 8,316 J/mol [7]
$A$ 3.34$\times$$10^{-3}$ s-1 [7]
[…]
$T_{0}$ 241.15 K [43]
$T_{b,0}$ 253.15 K [43]
$T_{b,\textrm{max}}$ 313.15 K [6]
$c_{s,0}$ 0.2059 kg water/kg solid [7]
$r$ 0.2 K/min [44]
$Q_{v}$ 0 W/m3
$H$ 2 cm [6]
[…]
$m$ 20
[…]
@@
Since the mechanistic model is a basis for state observer and control design, this section extensively validates the proposed model with three different datasets from the literature to ensure that our model provides an accurate prediction of the bound water concentration and product temperature. Table 2 lists the parameter values specific to the three case studies used for model validation; other parameters are based on the default values given in Table 1.
Table 2: Specific parameters for the model validation.
Case Parameter Value Unit Reference/Note
1 $E_{a}$ 5,000 J/mol Estimated from data
$A$ 7.1$\times$10-4 s-1 Estimated from data
$c_{s,0}$ 0.6415 kg water/kg solid [6]
2 $E_{a}$ 5,700 J/mol Estimated from data
$A$ 1$\times$10-3 s-1 Estimated from data
$c_{s,0}$ 0.6415 kg water/kg solid [13]
3 $E_{a}$ 5,920 J/mol [17]
$A$ 1.2$\times$$10^{-3}$ s-1 Estimated from data
@@ 3 $E_{a}$ 5,920 J/mol [17]
$T_{0}$ 264.09 K [17]
$T_{b,0}$ 264.09 K Assumed to be $T_{0}$
$T_{b,\textrm{max}}$ 312 K Estimated from data
$c_{s,0}$ 0.0603 kg water/kg solid [17]
$r$ 0.5 K/min Estimated from data
$H$ 0.0102 m Calculated from parameters in Ref. [17]
The first dataset, denoted Case 1, is obtained from the experimental data presented in Ref. [6], where the time profile of the total mass of residual water (bound water) during secondary drying was reported. Our mechanistic model can accurately predict the concentration of bound water, with the maximum error of about 0.01 kg water/kg solid (Fig. 3A). The concentration decreases exponentially following the linear driving force model.
The second dataset (Case 2) is the simulation result obtained from the high-fidelity model proposed by Ref. [13]; the model simulates simultaneous heat and mass transfer in two dimensions. Our model prediction agrees well the result obtained from the high-fidelity model (Fig. 3B), indicating that our simplified model, which simulates the system in 1D and omits mass transfer equations, is sufficiently accurate to be used for predicting the evolution of residual water during secondary drying, agreeing with the observation by Ref. [32].
The final dataset (Case 3) is obtained from Ref. [17], where the time profiles of both residual water and product temperature were reported. Our model can precisely simulate the product temperature (Fig. 3c). The model can also reasonably predict the concentration of bound water despite high uncertainty in the reported measurement.
Figure 3: (A) Comparison between the concentration of bound water predicted by our model and the experimental data obtained from [6]. The original data reported as the total mass of water are normalized by the total mass of solid. (B) Comparison between the concentration of bound water predicted by our model and the high-fidelity model by [13]. The original data reported as the total mass of water are normalized by the total mass of solid. (C) Comparison between our model prediction and the experimental data from [17] for both temperature and concentration.
@@
This section demonstrates the simulation-based observer design, with the default model parameters in Table 1. As introduced in Section 3.3, a practical design technique is to run a series of simulations for different values of the observer gains. Alternatively, the observer gains can be systematically selected from mathematical analysis of the observer (40)–(49) to produce a desired convergence time for the estimation error.
Figure 4 shows the convergence times for orders of magnitude ranges in observer gains; the convergence time can be estimated using (49) or directly calculated from simulation. Within this design space, the convergence time ranges from 0.7 to 9.3 h. For the process time scale of about 10 h, we select the observer gains $L_{T}=-1$$\times$$10^{-6}$ and $L_{c}=5$$\times$$10^{-7}$ to give a convergence time of about 2 h (dark blue region of Fig. 4), i.e., 5 times faster than the process dynamics.
Figure 4: Design space showing the times required for the estimated concentration to converge to the true value at different pairs of the observer gains $L_{T}$ and $L_{c}$. Values of $L_{T}$ and $L_{c}$ outside this design space could lead to severe oscillation or divergence.
With the selected observer gains, the estimated temperature quickly converges to the true value without any oscillation (Fig. 5A), with the estimation error decreasing to 0 at around 2 h (Fig. 5B). A similar behavior can be observed for the concentration (Fig. 6), indicating that the observer design is appropriate. Physically, when the estimated temperature is higher than the actual value, the observer term $\mathbf{L_{T}}(\hat{\mathbf{T}}-\mathbf{T})$ should provide negative feedback to the observer to reduce the temperature, so $L_{T}$ is negative. When the estimated temperature is higher than the actual value, it implies that the estimated desorption rate is too low. Therefore, the observer term $\mathbf{L_{c}}(\hat{\mathbf{T}}-\mathbf{T})$ should provide positive feedback to the observer to increase the estimated concentration, and so $L_{c}$ is positive.
Figure 5: Convergence of the estimated product temperature for the selected observer gains $L_{T}=-1$$\times$$10^{-6}$ and $L_{c}=5$$\times$$10^{-7}$. Panel A shows the evolution of the estimated average temperature compared to the true value. Panel B shows the spatiotemporal evolution of the estimation error. The initial estimated temperature is set to be higher than the actual value by about 10% to demonstrate convergence.
Figure 6: Convergence of the estimated concentration of bound water for the selected observer gains $L_{T}=-1$$\times$$10^{-6}$ and $L_{c}=5$$\times$$10^{-7}$. Panel A shows the evolution of the estimated average concentration compared to the true value. Panel B shows the spatiotemporal evolution of the estimation error. The initial estimated concentration is set to 0.0314 kg water/kg solid, the minimum value reported in the literature.
@@
Another important aspect of the observer is initialization. For the estimated temperature, it is logical to set the initial condition to be equal to the measured temperature. In Fig. 5, the initial estimated temperature is set to be higher than the measured value by 10% to demonstrate the convergence. The initial estimated concentration is, however, unknown due to no real-time concentration measurement. In Fig. 6, the observer is initialized with the minimum concentration (0.0314 kg water/kg solid) reported in the literature [6, 49] to demonstrate the convergence under the worst-case scenario. In practice, the initial estimated concentration could be set to some more realistic value specific to that system or experiment. This information should be obtained during the design and development phase, which could result in faster convergence. In the later sections of this work, the initial estimated concentration is set to 0.0314 kg water/kg solid and the initial estimated temperature is set to the measured value unless otherwise specified so that the observer performance is analyzed on the same basis.
In the later sections of this work, $L_{T}=-1$$\times$$10^{-6}$ and $L_{c}=5$$\times$$10^{-7}$ are used as the default observer gains. This exact same analysis and design procedure can be used for the modified observer, with the final observer gains selected to be $L_{T}=-5$$\times$$10^{-3}$ and $L_{c}=1$$\times$$10^{-4}$.
@@
Figure 7: (A) Convergence of the estimated concentration for different frequency factors $A$. The frequency factor is varied between 7.8$\times$10-5 s-1 [13] and 1.1$\times$10-4 s-1 [12] given that the rate constant is independent on temperature. (B) Convergence of the estimated concentration for different activation energies $E_{a}$. The activation energy is varied between 5,920 J/mol [17] and 13,146 J/mol [50]. (C) Convergence of the estimated concentration for different initial concentrations, $c_{s,0}$. The initial condition is varied between 0.0314 kg water/kg solid [4] and 0.6415 kg water/kg solid [6]. In all cases, the simulations are terminated when the estimated concentration drops below 0.01 kg water/kg solid. The default observer gains $L_{T}=-1$$\times$$10^{-6}$ and $L_{c}=5$$\times$$10^{-7}$ are used. Other parameters are kept at the default values.
@@
To simulate the noise, independent normally distributed noise of standard deviation given by $3\sigma=5^{\circ}$ [39] is added to the temperature profile given the default model parameters. The estimated concentration is severely polluted by the measurement noise for the default observer gain $L_{c}=5$$\times$$10^{-7}$ (Fig. 8A). Significant oscillation is observed, in which the observer does not give valuable information when the concentration is small (i.e., after about 6 h). To reduce the noise effect, the observer gain $L_{c}$ is reduced to $2$$\times$$10^{-7}$. In comparison to $L_{c}=5$$\times$$10^{-7}$, the convergence is achieved slightly slower, but the oscillation is much weaker. Reducing the gain $L_{c}$ further to $1$$\times$$10^{-7}$ almost removes the oscillation, but the convergence is also significantly slower.
The above information motivates the use of observer gain scheduling, which is a technique that varies the observer gains with time to ensure satisfactory performance at different operating points, especially for nonlinear systems. In this case, we can initialize the observer with a high gain to achieve fast convergence, and then gradually switch to a low gain after some time to reduce oscillation. Here the switching time is set to $t=4\tau$ following the definition of the convergence time defined in Section 3.3, where $\tau$ is the time constant derived in Section 3.3. This technique enables fast convergence while having low effects of measurement noise on the state estimate (Fig. 8B).
@@
Note that the above noise level (Fig. 8C) is significantly higher than that observed in actual experiments [39], and so the resulting design has a large safety margin to span the range of noise levels encountered in real temperature sensors. A lower safety margin, with faster convergence to the states, could be achieved by performing the analysis for a noise level set by experimental data for the specific temperature sensor used in the specific equipment.
Figure 8: (A) Convergence and oscillation behavior of the estimated concentration at different observer gains $L_{c}=5$$\times$$10^{-7}$, $2$$\times$$10^{-7}$, and $1$$\times$$10^{-7}$ while $L_{T}$ is fixed at $-1$$\times$$10^{-6}$ in all cases. (B) Gain scheduling technique where $L_{c}$ is set to $5$$\times$$10^{-7}$ for fast convergence at the beginning and then reduced to $1$$\times$$10^{-7}$ at $t=4\tau$ to reduce the effect of measurement noise. (C) Noisy temperature data obtained from adding independent normally distributed noise of standard deviation given by $3\sigma=5^{\circ}$ to the true state. Other parameters are kept at their default values.
@@
In the simulation-based observer design, the true and estimated states are simulated simultaneously and continuously. In real systems, measurement data are usually sampled and fed to a state observer in a discrete fashion, i.e., with a fixed sampling time. This section applies the proposed observer to estimate the concentration of bound water in four different experiments from the literature, denoted as Cases A (skim milk) [6], B (skim milk) [1], C (sucrose) [49], and D (mannitol) [17], with the sampling time of 10 s. Although our model is developed based on lyophilization of unit doses (vials), the experimental data used here include both bulk lyophilization (Cases A and B) and vial lyophilization (Cases C and D) to demonstrate the generalizability of the approach given that radial heat transfer is negligible compared to vertical heat transfer.
Table 3 lists the parameter values specific to the four case studies considered here. For Cases A and B, temperature data were not directly reported in the original studies [1, 6]; however, a mechanistic model and model parameters were given such that spatial temperature data could be simulated. The simulated temperature profile was used as the temperature measurement for our state observer, with the sampling time of 10 s. Additionally, the original residual moisture was reported as the total mass of water, so we normalized the values by the total mass of solid to give the unit of kg water/kg solid. For Cases C and D, the bottom temperatures were reported in the original studies [49, 17], but the sampling time is too large (about 30 min), resulting in poor convergence. Hence, based on the reported temperature, we used the model to reconstruct the the temperature profile with the sampling time of 10 s, which was then fed to our observer.
Table 3: Specific parameters for the four experimental systems.
Case Parameter Value Unit Reference/Note
1 $E_{a}$ 5,000 J/mol Estimated from data
$A$ 7.1$\times$10-4 s-1 Estimated from data
$c_{s,0}$ 0.6415 kg water/kg solid [6]
[…]
$E_{a}$ 5,300 J/mol Estimated from data
$A$ 4.5$\times$$10^{-4}$ s-1 Estimated from data
$c_{s,0}$ 0.1940 kg water/kg solid [1]
C $E_{a}$ 37,714 J/mol [49]
$A$ 277 s-1 [49]
[…]
$T_{0}$ 270.38 K [49]
$T_{b,0}$ 270.38 K Assumed to be $T_{0}$
$c_{s,0}$ 0.0314 kg water/kg solid [49]
$r$ 0.6 K/min Estimated from data
D $E_{a}$ 5,920 J/mol [17]
$A$ 1.2$\times$$10^{-3}$ s-1 Estimated from data
[…]
$T_{0}$ 264.09 K [17]
$T_{b,0}$ 264.09 K Assumed to be $T_{0}$
$T_{b,\textrm{max}}$ 312 K Estimated from data
$c_{s,0}$ 0.0603 kg water/kg solid [17]
$r$ 0.5 K/min Estimated from data
$H$ 0.0102 m Calculated from parameters in Ref. [17]